The rate of change of x is proportional to x. When t=0, x(0) =3, and when t=2, x=6. What js the value of x when t=4?


1
Expert's answer
2021-06-11T03:04:36-0400

"Question :-
\\
The \space rate \space of \space change \space of \space x \space is \space proportional \space to \space x.\\ When \\\space t=0, x(0) =3,\\ and \\when \\t=2, x(2)=6.\\
 What \space the \space value \space of \space x \space when \space t=4?

\\
-----------------------------------\\Solution:-
\\given\\
The \space rate \space of \space change \space of \space x \space is \space proportional \space to \space x.\\
hence \\
\implies \frac{dx}{dt} \propto x\\
\implies \frac{dx}{dt} = k. x\\
\implies \frac{dx}{x} = k.dt\\
integration \space both \space side \\
\implies log(x)=k.t+ c\\
\implies x=e^{k.t+ c}\\
now \space \\
use \space first\space condition
\\
When \\\space t=0, x(0) =3,\\ 
hence\\
\implies 3=e^{k.0+ c}\\
\implies 3=e^{c}\\
\implies c=log(3)\\
now \space \\
use \space second \space condition
\\
When \\\space t=2, x(2) =6,\\ 
hence\\
\implies 6=e^{k.2 +log(3)}\\
\implies6=e^{k.2 }.3\\
\implies 2=e^{k.2}\\
\implies log(2)={k.2}\\
\implies 0.3010={k.2}\\
\implies k=0.1505\\
hence \\
\implies x=e^{k.t+ c}\\and \space k=0.1505, c=log(3)\\
hence\\
\implies x=e^{0.1505.t}.3\\
now \space we \space find \space the \space value \space of \space x \space when \space t=4?\\
\implies x=e^{0.1505.*4}.3\\
\implies x=5.477"